OpenStudy (anonymous):

best answer receives fan and medal part a.) the area of a square is (16x^2-8x+1) square units. determine the length of each side of the square by factoring the areas expressions completely. show all work part b.) the area of a rectangle is (81x^2-4y^2)square units. determine the dimensions of the rectangle by factoring the area of the expression. show all work part c.) the volume of a rectangular box is (x^3-9x^2-4x+36) cubic units. determine the dimensions of the rectangular box by factoring the volume expression completely. show all work

OpenStudy (anonymous):

a.\[16 x^2-4x-4x+1=4x(4x-1)-1(4x-1)=\left( 4x-1 \right)(4x-1)=\left( 4x-1 \right)^2\] Length of each side=4x-1 b.\[a^2-b^2=(a+b)(a-b)\] c.\[x^2(x-9)-4(x-9)=(x-9)(x^2-4)\] make also the factors of \[x^2-4\]

OpenStudy (anonymous):

can you explain @surjithayer

OpenStudy (anonymous):

im kinda confused

OpenStudy (anonymous):

\[81 x^2-4y^2=(9x)^2 -(4y)^2=(9x+4y)(9x-4y)\]

OpenStudy (anonymous):

\[x^2-4=x^2-2^2=(x+2)(x-2)\]

OpenStudy (anonymous):

I mean can you explain using words, I don't really understand the picture equations

OpenStudy (anonymous):

a. \[area~ of~a~ \square= edge^2,edge=length~of~\square=4x-1\]

OpenStudy (anonymous):

b.area of a rectangle\[=length*width,length=x+2,width=x-2\]

OpenStudy (anonymous):

volume of rectangular box=length*width*height length=x+2,width=x-2,height=x-9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw